cookiemonster:

Determine the number of protons, electrons, and neutrons present in an atom of calcium (Ca). Explain how you determined your answer using complete sentences.

cookiemonster:

@Elsa213 @AnimeGhoul8863 @563blackghost @UsArmy3947

cookiemonster:

@Nnesha

Nnesha:

periodic table.

cookiemonster:

what??

cookiemonster:

oh, do you need the picture of the periodic table?

cookiemonster:

http://prntscr.com/jb1po9

Nnesha:

ah no i got plenty in my folder..

Nnesha:

Top top number represent the number of electrons https://prnt.sc/jb1q73

cookiemonster:

okay...

cookiemonster:

how about i tell you the answer i got and think it is and you fix it if it needs it :)

cookiemonster:

Part 1: Potassium and Rubidium Potassium and Rubidium are both Alkali Metals, as such is Sodium. They are alike in many ways, as is why they are categorized in the same group (Alkali Metals).  The properties of Alkali Metals are as follows: they have low melting and boiling points compared to most other metals.they are very soft and can be cut easily with a knife.they have low densities (lithium, sodium, and potassium will float on water) Part 2: Calcium has 20 protons. The reason why is because the atomic number is 20 (the one at the top). It has 20 neutrons. The reason why is because the atomic mass (the number on the bottom) is 40, ignoring electron mass. 40- 20 protons = 20 neutrons. It has 20 electrons. The reason why is because the number of protons and electrons are the same.

Nnesha:

"ignoring electron mass "??

Nnesha:

Mass number of atom = # of protons + # of neutrons

cookiemonster:

ok so how would you answer part b ?

Nnesha:

you did a good job!

cookiemonster:

oh so then why did you say, "ignoring electron mass "?? did i mess up there ? or do i take that part out or what?

Nnesha:

I was asking the explanation of that part.

cookiemonster:

ohhh okay, are you sure i didnt make any mistakes?

cookiemonster:

nvm i trust you. i think if you say i did good and didnt make a mistake i believe you :) i have some more multiple choice questions

Nnesha:

i will give you full point i'm not sure how strict your professor is. or for the middle sentence you can add one more sentence atomic mass = # of protons + # of neutrons therefore 40-20 = 20 neutrons

cookiemonster:

ok sounds great so it would be this: Calcium has 20 protons. The reason why is because the atomic number is 20 (the one at the top). It has 20 neutrons. The reason why is because the atomic mass (the number on the bottom) is 40, ignoring electron mass. atomic mass = # of protons + # of neutrons therefore 40-20 = 20 neutrons. 40- 20 protons = 20 neutrons. It has 20 electrons. The reason why is because the number of protons and electrons are the same.

Nnesha:

I have to go in few seconds so no promisez..

cookiemonster:

ok :) but does it look good ^^?

Nnesha:

repeated words

cookiemonster:

oh ok ill fix that... Calcium has 20 protons. The reason why is because the atomic number is 20 (the one at the top). It has 20 neutrons. The reason why is because the atomic mass (the number on the bottom) is 40, ignoring electron mass. atomic mass = # of protons + # of neutrons therefore 40-20 = 20 neutrons. It has 20 electrons. The reason why is because the number of protons and electrons are the same.

Nnesha:

Looks good.

cookiemonster:

great! next :)

Nnesha:

u can post it will see. like i said i gtg maybe someone else can help you.

cookiemonster:

What is a mixture called when it has its different components mixed unevenly within the substance? Heterogeneous mixture Homogeneous mixture Impure mixture Pure mixture